Calculate the energy values of the electron in second and third orbits of hydrogen atom
1
Expert's answer
2020-03-30T10:50:06-0400

for electron on the n-th orbit in hydrogen atom energy E is defined as:

E = -ER/n2

where ER = 13.6 eV is rydberg energy.

Energy of electron on the 2nd orbit is

E2 = -13.6/4 = -3.40 eV

And for the 3rd orbit:

E3 = -13.6/9 = 1.51 eV